# Conveyor-belt superconducting quantum computer

### v1

@@ II.2 Dynamical features
II.3 Information encoding and exchange operations
@@
and $|\Psi;{\rm PF}\rangle=|\Psi;{\rm FP}\rangle|_{|{\rm P}\rangle\leftrightarrow|{\rm F}\rangle}$. We notice that this requires $N$ to be an even number. When all qubits are initially in the ground state, the control $V_{\rm init}(t)$ can be used to initialize the system to a vector of the form $|\Psi;{\rm FP}\rangle$, see Fig. 1. This process brings all the associated qubits into the $|e\rangle$ state. As mentioned before, in our architecture, single and multi-qubit gates can then be performed on specific sites which host, or which are directly coupled to, crossed qubits elements (i.e. $Q_{1}$, $Q_{2}$ and $Q_{3}$). This implies that a fundamental prerequisite to perform quantum computation in our conveyor-belt architecture is the ability to coherently exchange the positions of the computational qubits, i.e. change the positions of the qubit states without modifying them. This is achieved trough a sequence of eight alternating global pulses, i.e. $\hat{\Pi}_{\rm exc}:=\hat{\Pi}_{B}\hat{\Pi}_{A^{\rm r}}\hat{\Pi}_{B}\hat{\Pi}_{A^{\rm r}}\hat{\Pi}_{B}\hat{\Pi}_{A^{\rm r}}\hat{\Pi}_{B}\hat{\Pi}_{A^{\rm r}}$ with $\hat{\Pi}_{A^{\rm r}}:=\hat{W}_{A^{\rm r}}(\pi,\bm{x})$ acting as a conditional-bit-flip only on the regular $A$-type qubits, and $\hat{\Pi}_{B}:=\hat{W}_{B}(\pi,\bm{x};\pi,\bm{x})$ acting as a conditional bit-flip on all $B$-type qubits, including the crossed one. As shown in the supplemental material supplemental , when acting on a ${\rm FP}$ (resp. ${\rm PF}$) well-formed state $|\Psi;{\rm FP}\rangle$ ($|\Psi;{\rm PF}\rangle$), the transformation $\hat{\Pi}_{\rm exc}$ runs in parallel $N$ two-qubit swap gates $\hat{U}^{{\rm swap}}_{Q_{j}Q_{j+1}}$ on the pairs $\{Q_{1},Q_{2}\}$, $\{Q_{3},Q_{4}\}$, $\cdots$, $\{Q_{N-1},Q_{N}\}$, (resp. $\{Q_{2},Q_{3}\}$, $\{Q_{4},Q_{5}\}$, $\cdots$, $\{Q_{N},Q_{1}\}$) while exchanging the ferromagnetic and paramagnetic phases of the intermediate sectors producing a ${\rm PF}$ (resp. ${\rm FP}$) output configuration—see Fig. 2. The resulting motion of the computational qubits acquires an interesting feature. Indeed one has that $\hat{\Pi}_{\rm exc}|\Psi;{\rm FP}\rangle=|\Psi^{(1)}_{{\circlearrowright}_{o},{\circlearrowleft}_{e}};{\rm PF}\rangle$, and $\hat{\Pi}_{\rm exc}|\Psi;{\rm PF}\rangle=|\Psi^{(1)}_{{\circlearrowleft}_{o},{\circlearrowright}_{e}};{\rm FP}\rangle$, with the vector $|\Psi^{(\ell)}_{{\circlearrowright}_{o},{\circlearrowleft}_{e}}\rangle$ (resp. $|\Psi^{(\ell)}_{{\circlearrowleft}_{o},{\circlearrowright}_{e}}\rangle$) obtained by applying to $|\Psi\rangle$ an $\ell$-step clock-wise rotation of the internal states of the sites $Q_{j}$ with odd (even) index $j$ and, at the same time, an $\ell$-step anti-clock-wise rotation of the internal states of the sites $Q_{j}$ with even (odd) index $j$. Notice that since the direction of the rotations depends on the initial location of the ferromagnetic and paramagnetic phases supplemental , multiple applications of $\hat{\Pi}_{\rm exc}$ pulses do not cancel out. For instance, using $\ell$ times the transformation on the input state $|\Psi;{\rm FP}\rangle$ leads to $\hat{\Pi}^{\ell}_{\rm exc}|\Psi;{\rm FP}\rangle=|\Psi^{(\ell)}_{{\circlearrowright}_{o},{\circlearrowleft}_{e}};{\rm FP}\rangle$ for $\ell$ even, and $\hat{\Pi}^{\ell}_{\rm exc}|\Psi;{\rm FP}\rangle=|\Psi^{(\ell)}_{{\circlearrowright}_{o},{\circlearrowleft}_{e}};{\rm PF}\rangle$ for $\ell$ odd (of course for $\ell=N$ the system goes back to the initial configuration).
@@
Exploiting this feature, we can move any IC qubit in any other IC qubit location. Suppose, now, that the aim is to apply the single qubit transformation ^$\hat{\mathbb{R}}(\theta,\bm{n})$ to the computational qubit located (say) in the $Q_{j}$ site of the loop. Given our previous explanations, this task is straightforward. We simply need to use the transformation $\hat{\Pi}^{\ell}_{\rm exc}$ with $\ell$ such that the state of $Q_{j}$ rotates into the position $Q_{2}$ which is hosting the $B$-type crossed qubit. Once there, we apply the pulse $\hat{U}_{B^{\times}}:=\hat{W}_{B^{\times}}(\theta,\bm{n})$ and (in case we need it), reverse the exchange operation via the sequence $\hat{\Pi}^{N-\ell}_{\rm exc}$ (or using $\ell$-times the inverse of $\hat{\Pi}^{\ell}_{\rm exc}$ defined in the caption of Fig. 2). This demonstrates that we are capable of performing any single-qubit unitary operation on any one of the computational qubits. In the next Section, we describe how a three-qubit Toffoli gate can be implemented, thereby achieving a universal gate set.
II.4 One-shot Toffoli gate
@@
In common QC architectures, the Toffoli gate is performed using two-qubit gates. Here, we implement a three-qubit one-shot Toffoli gate by using a total number of four qubits: one $A$-type crossed qubit located inside the loop, which acts as a mediator, and the three $B$-type computational qubits located in the IC sites $Q_{1}$, $Q_{2}$ and $Q_{3}$. A similar setup has already been proposed by Rasmussen et alRasmussen202_toffoli as an isolated multi-qubit gate. The most notable feature of our Toffoli gate is that it is naturally embedded into our conveyor-belt QC architecture. Here, the natural way to implement a Toffoli gate is to decompose it into a controlled-controlled-Z (CCZ) gate combined with two single-qubit Hadamard gates. Since we have already discussed how to perform single-qubit gates, we simply need to explain now how to perform a CCZ gate. The fundamental observation here is that the unitary $\hat{\mathcal{Z}}_{A^{\times}}:=\hat{W}_{A^{\times}}(2\pi,\bm{n})$ induces a $(-1)$ phase factor on the state of the system if and only if the three $B$-type qubits connected to it are all in the ground state $|g\rangle$ which, apart from a global NOT, is exactly a CCZ gate applied on the three $B$-type qubits. Specifically, the transformation $\hat{\mathcal{Z}}_{A^{\times}}$ is obtained from (1) for $\chi=A$ and by setting $\theta^{\prime}=0$ and $\theta^{\prime\prime}=2\pi$, independently from the choices of $\bm{n}^{\prime}$ and $\bm{n}^{\prime\prime}$. It corresponds to $\hat{\mathcal{Z}}_{A^{\times}}=\hat{Q}_{\langle A^{\times}\rangle}-\hat{P}_{\langle A^{\times}\rangle}$, where, since the $A$-type crossed qubit has three connections, the projector $\hat{P}_{\langle A^{\times}\rangle}$ can be written as $\hat{P}_{\langle A^{\times}\rangle}:=|ggg\rangle\langle ggg|$, while the complementary projector is $\hat{Q}_{\langle A^{\times}\rangle}:=|eee\rangle\langle eee|+|eeg\rangle\langle eeg|+\cdots+|gge\rangle\langle gge|$. Accordingly, setting $\bm{n}=(1,0,1)/\sqrt{2}$, the transformation
@@
We conclude this Section by remarking that this whole discussion can be generalized to an arbitrary number of qubits connected to the central $A$-type crossed qubit. It is therefore possible, at least in principle, to realize $N$-qubit Toffoli gates in our globally-controlled conveyor-belt architectures.
II.5 Universality
@@
In the previous Sections, we have demonstrated how, by using the global controls $V_{A}(t)$ and $V_{B}(t)$, the architecture in Fig. 1 allows us to implement any single-qubit operation on each site that encodes the logical information, as well as to perform the exchange gate $\hat{\Pi}_{\rm exc}$ and the Toffoli gate $\hat{T}_{13\rightarrow 2}$. To prove that this is sufficient to guarantee universal QC in the model, we now show that, using these gates, it is possible to induce arbitrary swap gates $\hat{U}^{{\rm swap}}_{Q_{j}Q_{j^{\prime}}}$ between any two IC sites. These will then be used to convert $\hat{T}_{13\rightarrow 2}$ into Toffoli gates that couple all possible triples formed by computational qubits. It should first be noted that the Toffoli gate, when combined with single-qubit operations, can be employed to implement local swaps between any two qubits that are simultaneously connected to the central $A$-type crossed qubit Nielsen2010 . Starting from $\hat{T}_{13\rightarrow 2}$, we can generate swap gates between each of the pairs $\{Q_{1},Q_{2}\}$, $\{Q_{2},Q_{3}\}$, and $\{Q_{3},Q_{1}\}$ that we represent as edges of a $N$ vertex graph formed by the computational qubits of the system, see Fig. 2(b). Next, we apply $\hat{\Pi}_{\rm exc}$ to rotate the computational qubits, e.g. inducing the mapping $(Q_{1},Q_{2},Q_{3})\leftrightarrow(Q_{2},Q_{1},Q_{4})$, and repeat the whole procedure obtaining two new swaps among the couple $\{Q_{1},Q_{4}\}$ and $\{Q_{2},Q_{4}\}$ which allows us to draw two new edges in the graph. Note also that, by combining these swaps, we can also induce an extra new swap between $\{Q_{3},Q_{4}\}$ (e.g. $\hat{U}^{{\rm swap}}_{Q_{3}Q_{4}}=\hat{U}^{{\rm swap}}_{Q_{1}Q_{3}}\hat{U}^{{\rm swap}}_{Q_{1}Q_{4}}\hat{U}^{{\rm swap}}_{Q_{1}Q_{3}}$), making the subgraph formed by the sites $Q_{1}$, $Q_{2}$, $Q_{3}$ and $Q_{4}$ fully connected. Proceeding along this way, one can show by recursion that all the other vertices can also be included in a fully connected graph, meaning that we can generate all possible swaps $\hat{U}^{{\rm swap}}_{Q_{j}Q_{j^{\prime}}}$, hence proving the thesis supplemental .
III Discussion
@@
The model we proposed represents a tremendous improvement in terms of scalability compared to those discussed in Refs. menta2024globally ; cesa2023universal where one needs $\mathcal{O}(N^{2})$ physical qubits in order to have a universal quantum computer with $N$ computational qubits. Indeed, in the present work, we have presented a quantum processing unit where the scaling is $\mathcal{O}(N)$. Globally-driven schemes with linear scaling have already been proposed  Lloyd_1993 ; benjamin_2000 ; benjamin_2001 ; benjamin_2003 ; benjamin_2004 . In contrast to our proposal, however, these ones remain rather abstract, since they do not provide precise recipes on how to actually implement the proposed global schemes. Additionally, there are many elements of difference in how the computation is performed. Among these, the processor described in the present work does not rely on the dynamical tuning of the system’s parameters, such as level spacings and couplings. Also, we do not use a redundant encoding of the computational qubits. A single, two-level system contains the information of a single computational qubit. This greatly simplifies in our humble opinion the extension of our conveyor-belt architecture to the level that is needed for fault-tolerant QC.
@@
All considerations regarding the physical requirements (e.g. values of the Rabi frequencies, level spacings, ZZ coupling strengths, etc.) for the implementation of our setup remain consistent with those presented in Ref. menta2024globally , since the physical building blocks involved are identical. Another novelty introduced by the setup in Fig. 1, as compared to the one proposed in Ref. menta2024globally , is the ability of the former to perform one-shot Toffoli gates Zahedinejad2015 ; Rasmussen202_toffoli ; baker2022 . This could in principle improve the execution time as well as the fidelity of many quantum algorithms. From a practical standpoint, the fabrication of a single large loop, containing many qubits, and the design of global control lines that simultaneously connect all the qubits may prove to be an inefficient and error-prone process. Integrating a hybrid control with different global sources for different subsets of qubits, fabricated on different chiplets, could prove advantageous.
A major challenge of our architecture, from the experimental point of view, arises from the fact that spatial inhomogeneities in the fabrication of superconducting qubits and in the always-on ZZ couplings throughout the conveyor belt can influence the fidelity of the necessary logical operations. While a thorough disorder and noise analysis is well beyond the scope of this work, it is important to note that systematic fabrication errors can be managed as conventional dynamical errors within error correction frameworks.

### v2

@@ III.1 Effective blockade dynamics
[…]
$\hat{\mathbb{R}}_{i}(\theta,\bm{n}_{\perp})=\exp\!\left[-i\frac{\theta}{2}\bm{n}_{\perp}\!\cdot\!\vec{\sigma}^{(i)}\right]$ (4)
is the single-qubit rotation induced in the absence of interactions, with angle $\theta=\Omega_{\chi}\tau$ and axis $\bm{n}_{\perp}=(\cos\phi_{\chi},\sin\phi_{\chi},0)$.
III.2 Global control unitaries
Due to the geometry of the device, qubits of the same species $\chi$ split into two subsets defined as regular ($\chi^{\rm r}$) and crossed ($\chi^{\times}$). As mentioned, these qubits are characterized by different Rabi frequencies. It can be shown that this inhomogeneity is sufficient to achieve independent control of the two subsets, even though the pulses acting on the qubits are global menta2024globally .
Due to the global nature of the drive, the evolution induced by a single pulse can be written compactly as
$\hat{U}_{\chi}=\hat{W}_{\chi}(\theta,\bm{n}_{\perp};2\theta,\bm{n}_{\perp}),$ (5)
where we define the global control unitaries
$\hat{W}_{\chi}(\theta^{\prime},\bm{n}^{\prime};\theta^{\prime\prime},\bm{n}^{\prime\prime})=\hat{W}_{\chi^{\rm r}}(\theta^{\prime},\bm{n}^{\prime})\hat{W}_{\chi^{\times}}(\theta^{\prime\prime},\bm{n}^{\prime\prime}).$ (6)
Recalling the above mentioned result, for $\xi\in\{\chi^{\rm r},\chi^{\times}\}$, we have
[…]
with $\hat{\mathbb{R}}_{i}(\theta,\bm{n})=\exp[-i(\theta/2)\bm{n}\cdot\vec{\sigma}_{i}]$ a single-qubit $SU(2)$ rotation Nielsen2010 . Depending on the state of the neighboring sites, each qubit either undergoes the rotation $\hat{\mathbb{R}}_{i}(\theta,\bm{n})$ or remains unchanged.
When only one of the parameters $\theta^{\prime}$ or $\theta^{\prime\prime}$ in Eq. (6) is nonzero, the corresponding transformation acts selectively on either the regular or crossed subset of qubits. Although the elementary pulses in Eq. (5) produce correlated rotations on the two subsets, suitable pulse sequences allow one to synthesize arbitrary rotations on $\chi^{\rm r}$ and $\chi^{\times}$ independently, and hence to generate generic unitaries of the form $\hat{W}_{\chi}(\theta^{\prime},\bm{n}^{\prime};\theta^{\prime\prime},\bm{n}^{\prime\prime})$. The explicit constructions establishing the universality of such control are summarized in Appendix A and detailed in Ref. menta2024globally . A generalization of this result to the case of more than two subsets is proven in Ref. menta2025building .
IV Information encoding and exchange operations
@@
[…]
@@
Exploiting this feature, we can move any IC qubit in any other IC qubit location. Suppose, now, that the aim is to apply the single qubit transformation $\hat{\mathbb{R}}(\theta,\bm{n})$ to the computational qubit located (say) in the $Q_{j}$ site of the loop. Given our previous explanations, this task is straightforward. We simply need to use the transformation $\hat{\Pi}^{\ell}_{\rm exc}$ with $\ell$ such that the state of $Q_{j}$ rotates into the position $Q_{2}$ which is hosting the $B$-type crossed qubit. Once there, we apply the pulse $\hat{U}_{B^{\times}}:=\hat{W}_{B^{\times}}(\theta,\bm{n})$ and (in case we need it), reverse the exchange operation via the sequence $\hat{\Pi}^{N-\ell}_{\rm exc}$ (or using $\ell$-times the inverse of $\hat{\Pi}^{\ell}_{\rm exc}$ defined in the caption of Fig. 2). This demonstrates that we are capable of performing any single-qubit unitary operation on any one of the computational qubits. In the next paragraph, we describe how a three-qubit Toffoli gate can be implemented, thereby achieving a universal gate set.
V One-shot Toffoli gate
@@
In common QC architectures, the Toffoli gate is performed using two-qubit gates. Here, we implement a three-qubit one-shot Toffoli gate by using a total number of four qubits: one $A$-type crossed qubit located inside the loop, which acts as a mediator, and the three $B$-type computational qubits located in the IC sites $Q_{1}$, $Q_{2}$ and $Q_{3}$. A similar setup has already been proposed by Rasmussen et alRasmussen202_toffoli as an isolated multi-qubit gate. The most notable feature of our Toffoli gate is that it is naturally embedded into our conveyor-belt QC architecture. Here, the natural way to implement a Toffoli gate is to decompose it into a controlled-controlled-Z (CCZ) gate combined with two single-qubit Hadamard gates. Since we have already discussed how to perform single-qubit gates, we simply need to explain now how to perform a CCZ gate. The fundamental observation here is that the unitary $\hat{\mathcal{Z}}_{A^{\times}}:=\hat{W}_{A^{\times}}(2\pi,\bm{n})$ induces a (1)(-1) phase factor on the state of the system if and only if the three $B$-type qubits connected to it are all in the ground state $|g\rangle$ which, apart from a global NOT, is exactly a CCZ gate applied on the three $B$-type qubits. Specifically, the transformation $\hat{\mathcal{Z}}_{A^{\times}}$ is obtained from (6) for $\chi=A$ and by setting $\theta^{\prime}=0$ and $\theta^{\prime\prime}=2\pi$, independently from the choices of $\bm{n}^{\prime}$ and $\bm{n}^{\prime\prime}$. It corresponds to $\hat{\mathcal{Z}}_{A^{\times}}=\hat{Q}_{\langle A^{\times}\rangle}-\hat{P}_{\langle A^{\times}\rangle}$, where, since the $A$-type crossed qubit has three connections, the projector $\hat{P}_{\langle A^{\times}\rangle}$ can be written as $\hat{P}_{\langle A^{\times}\rangle}:=|ggg\rangle\langle ggg|$, while the complementary projector is $\hat{Q}_{\langle A^{\times}\rangle}:=|eee\rangle\langle eee|+|eeg\rangle\langle eeg|+\cdots+|gge\rangle\langle gge|$. Accordingly, setting $\bm{n}=(1,0,1)/\sqrt{2}$, the transformation
@@
We conclude this paragraph by remarking that this whole discussion can be generalized to an arbitrary number of qubits connected to the central $A$-type crossed qubit. It is therefore possible, at least in principle, to realize $N$-qubit Toffoli gates in our globally-controlled conveyor-belt architectures.
VI Universality
@@
In the previous paragraphs, we have demonstrated how, by using the global controls $V_{A}(t)$ and $V_{B}(t)$, the architecture in Fig. 1 allows us to implement any single-qubit operation on each site that encodes the logical information, as well as to perform the exchange gate $\hat{\Pi}_{\rm exc}$ and the Toffoli gate $\hat{T}_{13\rightarrow 2}$. To prove that this is sufficient to guarantee universal QC in the model, we now show that, using these gates, it is possible to induce arbitrary swap gates $\hat{U}^{{\rm swap}}_{Q_{j}Q_{j^{\prime}}}$ between any two IC sites. These will then be used to convert $\hat{T}_{13\rightarrow 2}$ into Toffoli gates that couple all possible triples formed by computational qubits. It should first be noted that the Toffoli gate, when combined with single-qubit operations, can be employed to implement local swaps between any two qubits that are simultaneously connected to the central $A$-type crossed qubit Nielsen2010 . Starting from $\hat{T}_{13\rightarrow 2}$, we can generate swap gates between each of the pairs $\{Q_{1},Q_{2}\}$, $\{Q_{2},Q_{3}\}$, and $\{Q_{3},Q_{1}\}$ that we represent as edges of a $N$ vertex graph formed by the computational qubits of the system, see Fig. 2(b). Next, we apply $\hat{\Pi}_{\rm exc}$ to rotate the computational qubits, e.g. inducing the mapping $(Q_{1},Q_{2},Q_{3})\leftrightarrow(Q_{2},Q_{1},Q_{4})$, and repeat the whole procedure obtaining two new swaps among the couple $\{Q_{1},Q_{4}\}$ and $\{Q_{2},Q_{4}\}$ which allows us to draw two new edges in the graph. Note also that, by combining these swaps, we can also induce an extra new swap between $\{Q_{3},Q_{4}\}$ (e.g. $\hat{U}^{{\rm swap}}_{Q_{3}Q_{4}}=\hat{U}^{{\rm swap}}_{Q_{1}Q_{3}}\hat{U}^{{\rm swap}}_{Q_{1}Q_{4}}\hat{U}^{{\rm swap}}_{Q_{1}Q_{3}}$), making the subgraph formed by the sites $Q_{1}$, $Q_{2}$, $Q_{3}$ and $Q_{4}$ fully connected. Proceeding along this way, one can show by recursion that all the other vertices can also be included in a fully connected graph, meaning that we can generate all possible swaps $\hat{U}^{{\rm swap}}_{Q_{j}Q_{j^{\prime}}}$, hence proving the thesis (see Appendix C for details).
VII Discussion
@@
The model we proposed represents a clear improvement in terms of scalability compared to those discussed in Refs. menta2024globally ; cesa2023universal where $\mathcal{O}(N^{2})$ physical qubits are needed in order to have a universal quantum computer with $N$ computational qubits. Indeed, in the present work, we have presented a quantum processing unit where the scaling is $\mathcal{O}(N)$. Globally-driven schemes with linear scaling have already been proposed  Lloyd_1993 ; benjamin_2000 ; benjamin_2001 ; benjamin_2003 ; benjamin_2004 . In contrast to our proposal, however, these ones remain rather abstract, since they do not provide precise recipes on how to actually implement the proposed global schemes. Additionally, there are many elements of difference in how the computation is performed. Among these, the processor described in the present work does not rely on the dynamical tuning of the system’s parameters, such as level spacings and couplings. Also, we do not use a redundant encoding of the computational qubits. A single, two-level system contains the information of a single computational qubit, greatly reducing the complexity of the architecture.
@@
Another novelty introduced by the setup in Fig. 1 is the ability to perform one-shot Toffoli gates Zahedinejad2015 ; Rasmussen202_toffoli ; baker2022 . This could, in principle, improve the execution time and the fidelity of many quantum algorithms. The proposed architecture turns out to be very versatile. Indeed, universality could be achieved by substituting the Toffoli gate with two entangling gates as discussed in Appendix E. This versatility, together with the potential modularity of this proposal might be useful. From a practical standpoint, the fabrication of a single large loop containing many qubits and the design of global control lines that simultaneously connect all the qubits may prove to be an inefficient and error-prone process. Integrating a hybrid control with different global sources for different subsets of qubits, fabricated on different chiplets, could prove advantageous.
It is crucial to point out that a major experimental challenge of our architecture arises from spatial inhomogeneities in the fabrication of superconducting qubits and in the always-on ZZ couplings along the conveyor belt, which can affect the fidelity of the required logical operations. However, it has been recently demonstrated Aiudi2025 that by employing globally optimized control pulses, it is possible to overcome such disorder and achieve high fidelities. We stress that all considerations regarding the physical requirements (e.g. values of the Rabi frequencies, level spacings, ZZ coupling strengths, etc.) for the implementation of our setup remain consistent with those presented in Refs. menta2024globally ; Aiudi2025 , since the physical building blocks involved are identical.